OpenStudy (idealist10):

Solve for y: (y+e^x)/(y+2e^x)=Ce^x.

OpenStudy (phi):

first step is multply both sides by (y+2e^x) can you do that ?

OpenStudy (idealist10):

Yes, I got y+e^x=Ce^x(y+2e^x)

OpenStudy (phi):

next, "distribute" the C e^x on the right hand side i.e. multiply it times each term inside the parens

OpenStudy (idealist10):

I got y+e^x=yCe^x+2Ce^2x

OpenStudy (idealist10):

Now what?

OpenStudy (phi):

yes add -y to both sides add -2Ce^2x to both sides (i.e. get all the terms with y on one side, and the other terms on the other side)

OpenStudy (idealist10):

So I got e^x+2Ce^2x=yCe^x+y

OpenStudy (phi):

you did something wrong. if (for example) you add -y to both sides: y+e^x=yCe^x+2Ce^2x y - y + e^x = yCe^x -y + 2Ce^2x which simplifies to e^x = yCe^x -y + 2Ce^2x notice you have -y on the left side

OpenStudy (idealist10):

Now what?

OpenStudy (phi):

what do you have so far?

OpenStudy (idealist10):

e^x=yCe^x-y+2Ce^2x

OpenStudy (phi):

add -2Ce^2x to both sides (this makes all the terms on the left have a "y" in them)

OpenStudy (idealist10):

So I got e^x-2Ce^2x=yCe^x-y

OpenStudy (phi):

yes, much better. notice you have a y in both terms on the left side. any ideas how to "get at it" ?

OpenStudy (idealist10):

factor?

OpenStudy (phi):

yes

OpenStudy (idealist10):

Let me try

OpenStudy (phi):

it's like "undoing" distribute.

OpenStudy (idealist10):

Thank you so much, I got the right answer.

OpenStudy (phi):

yw